OpenStudy (anonymous):

The Lewis and Clark Expedition reported that it rained 94 days, which is 89% of their days in Oregon. How many days did the Lewis and Clark Expedition spend in Oregon? Round to the nearest whole day.

OpenStudy (anonymous):

This is a great problem for proportions. We know that 94 days is 89% of the total days in Oregon (super rainy, what a bummer). That means that 94 out of some number of days, let's call it D, is 89 percent. So 94/D=89/100. Now this simplifies to 94*100=89*D (cross multiplication). To solve, we get (94*100)/89=D. So D=105.62, or 106 days rounded to the nearest whole day. HTH!

OpenStudy (anonymous):

= 100 * 94/ 89 = 105.6 ~ 106 days

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

Any question?